OpenStudy (anonymous):

quick and easy question about circles (x+2)^2(y+2)^2 =121/4.......the center would be (-2,-2) or (2,2)

OpenStudy (anonymous):

11/2 is radius right?

OpenStudy (anonymous):

the center would be -2,-2 because you have to set x+2 and y+2 equal to 0 x+2=0 x=-2 y+2=0 y=-2

OpenStudy (anonymous):

the radius should be 11

OpenStudy (anonymous):

11/2

OpenStudy (anonymous):

but 121/4 squared is 11/2

OpenStudy (anonymous):

yep

OpenStudy (asnaseer):

the standard equation of a circle is given by:\[(x-h)^2+(y-k)^2=r^2\]so your radius is correct.

OpenStudy (anonymous):

oh my bad didnt notice the 121/4 so yeah its 11/2

OpenStudy (asnaseer):

(h, k) is the center of the circle, so in your case it would be at (-2, -2)

OpenStudy (anonymous):

asnaseer can you help me

OpenStudy (asnaseer):

if you want to learn more this might be useful: http://www.mathwarehouse.com/geometry/circle/equation-of-a-circle.php

OpenStudy (anonymous):

yea thanks i was just making sure my answers were correct

OpenStudy (asnaseer):

yw :)